# The Fefferman metric for twistor CR manifolds and conformal geodesics in dimension three

### v3

@@ 4.1. An adapted frame on $\widehat{\mathcal{C}}$
$\eta:=i\zeta\times\overline{\zeta}=2\,\mathrm{Re}\,\zeta\times\mathrm{Im}\,\zeta.$
@@
$\eta^{j}=i\varepsilon^{jkl}\zeta_{k}\overline{\zeta_{l}}.$
@@
$g(\eta,\eta)=1,\quad g(\eta,\zeta)=g(\eta,\overline{\zeta})=0,$
$\eta\times\overline{\zeta}=i\overline{\zeta},\quad\zeta\times\eta=i\zeta.$
@@
Since the conditions $g(\zeta,\zeta)=0,\ g(\zeta,\overline{\zeta})=1$ are preserved by the parallel transport of $\zeta$, we can define the horizontal lifts of tangent vectors on $\Sigma$ to $\widehat{\mathcal{C}}$. Let $\Gamma_{ij}{}^{k}$ be the Christoffel symbols of the Levi-Civita connection of $g$. Then, the horizontal lift of the tangent vector $\partial/\partial x^{i}$ is given by
@@
$\delta\zeta_{i}:=d\zeta_{i}-\Gamma_{ki}{}^{j}\zeta_{j}dx^{k},\quad\delta\overline{\zeta}_{i}:=d\overline{\zeta}_{i}-\Gamma_{ki}{}^{j}\overline{\zeta}_{j}dx^{k},$
@@
Using these notations, we define a frame $(\tilde{T},\tilde{Z}_{1},\tilde{Z}_{2},\tilde{Z}_{\overline{1}},\tilde{Z}_{\overline{2}},K)$ for $\mathbb{C}T\widehat{\mathcal{C}}$ and its dual coframe $(\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3})$ by
$\tilde{T}:=\eta^{i}\frac{\delta}{\delta x^{i}},\quad\tilde{Z}_{1}:=\overline{\zeta^{i}}\frac{\delta}{\delta x^{i}},\quad\tilde{Z}_{2}:=-i\eta_{j}\frac{\partial}{\partial\zeta_{j}},\quad K:=i\left(\zeta_{j}\frac{\partial}{\partial\zeta_{j}}-\overline{\zeta}_{j}\frac{\partial}{\partial\overline{\zeta}_{j}}\right)$
@@
$\theta:=\eta_{i}dx^{i},\quad\tilde{\theta}^{1}:=\zeta_{i}dx^{i},\quad\tilde{\theta}^{2}:=i\eta^{j}\delta\zeta_{j},\quad\tilde{\theta}^{3}:=-\frac{i}{2}(\overline{\zeta^{j}}\delta\zeta_{j}-\zeta^{j}\delta\overline{\zeta}_{j})=-i\overline{\zeta^{j}}\delta\zeta_{j}.$
@@
We shall prove that this indeed gives a (co)frame which is ‘adapted’ to the CR structure on $M$:
@@
(1) The $1$-from $\theta$ descends to a contact form on $M$.
(2) The $1$-forms $\tilde{\theta}^{1},\tilde{\theta}^{2}$ are horizontal with respect to $\widehat{\mathcal{C}}\to M$, and for any local section $s\colon M\to\widehat{\mathcal{C}}$, the set of $1$-forms $(\theta,\theta^{1}=s^{*}\tilde{\theta}^{1},\theta^{2}=s^{*}\tilde{\theta}^{2},\theta^{\overline{1}}=\overline{\theta^{1}},\theta^{\overline{2}}=\overline{\theta^{2}})$ gives an admissible coframe on $M$.
@@
(1) Since each coefficient $\eta_{i}$ is a $\mathrm{U}(1)$-invariant function, $\theta$ descends to a $1$-form on $M$. Moreover, by Proposition 3.3, $\theta$ annihilates $T^{1,0}M\oplus T^{0,1}M$ hence is a contact form on $M$. We note that the projection $T$ of $\tilde{T}$ to $M$ is well-defined since $\tilde{T}$ is $\mathrm{U}(1)$-invariant, and it gives the Reeb vector field of $\theta$. In fact, $T$ is represented as $T=\eta^{i}(\partial/\partial x^{i})$ at $p$, and since $d\varepsilon_{ijk}(x_{0})=0$, $dg_{ij}(x_{0})=0$, we have $d\eta_{i}(T)=0$, $\eta^{i}d\eta_{i}=0$ at $p$ and hence $\theta(T)=\eta_{i}\eta^{i}=1,\ T\lrcorner\,d\theta=T\lrcorner\,(d\eta_{i}\wedge dx^{i})=0.$
(2) Since $\tilde{\theta}^{1}(K)=\tilde{\theta}^{2}(K)=0$, these forms are horizontal. By Proposition 3.3, we see that $\theta^{1}$ and $\theta^{2}$ annihilate $T^{0,1}M$. Moreover, $\tilde{\theta}^{1}(\tilde{T})=\tilde{\theta}^{2}(\tilde{T})=0$ implies $\theta^{1}(T)=\theta^{2}(T)=0$. Thus, $(\theta,\theta^{1},\theta^{2},\theta^{\overline{1}},\theta^{\overline{2}})$ is an admissible coframe. ∎
@@
$\delta_{\lambda}^{*}\tilde{\theta}^{1}=\lambda\tilde{\theta}^{1},\quad\delta_{\lambda}^{*}\tilde{\theta}^{2}=\lambda\tilde{\theta}^{2}.$
Thus, $(\tilde{\theta}^{1},\tilde{\theta}^{2})$ can be considered as a ‘twisted’ or ‘weighted’ coframe for $T^{1,0}M$.
We also note that if we replace $g$ with $\widehat{g}=e^{2\Upsilon}g\ (\Upsilon\in C^{\infty}(\Sigma))$, we have $\widehat{\eta}^{\,i}=e^{-\Upsilon}\eta^{i}$, $\widehat{\eta}_{i}=e^{\Upsilon}\eta_{i}$ and hence the associated contact form is rescaled as $\widehat{\theta}=e^{\Upsilon}\theta$.
4.2. The differentials of the adapted coframe
@@
(4.1) $dx^{j}=\eta^{j}\theta+\overline{\zeta^{j}}\tilde{\theta}^{1}+\zeta^{j}\tilde{\theta}^{\overline{1}},\quad d\zeta_{j}=-i(\eta_{j}\tilde{\theta}^{2}-\zeta_{j}\tilde{\theta}^{3}),\quad d\eta^{j}=i(\overline{\zeta^{j}}\tilde{\theta}^{2}-\zeta^{j}\tilde{\theta}^{\overline{2}})$
@@
$dx^{j}(\tilde{T})=\eta^{j},\quad dx^{j}(\tilde{Z}_{1})=\overline{\zeta^{j}},\quad dx^{j}(\tilde{Z}_{2})=dx^{j}(K)=0,$
@@
$d\zeta_{j}(\tilde{T})=d\zeta_{j}(\tilde{Z}_{1})=d\zeta_{j}(\tilde{Z}_{\overline{1}})=d\zeta_{j}(\tilde{Z}_{\overline{2}})=0,\quad d\zeta_{j}(\tilde{Z}_{2})=-i\eta_{j},\quad d\zeta_{j}(K)=i\zeta_{j},$
and obtain the formula for $d\zeta_{j}$. Finally, using $d\varepsilon_{ijk}(x_{0})=0$, we compute as
$d\eta=id\zeta\times\overline{\zeta}+i\zeta\times d\overline{\zeta}=(\eta\tilde{\theta}^{2}-\zeta\tilde{\theta}^{3})\times\overline{\zeta}-\zeta\times(\eta\tilde{\theta}^{\overline{2}}-\overline{\zeta}\tilde{\theta}^{3})=i\overline{\zeta}\tilde{\theta}^{2}-i\zeta\tilde{\theta}^{\overline{2}}$
@@
(4.2) $d\theta$ $=i(\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{2}}+\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{1}}),$
$d\tilde{\theta}^{1}$ $=i\theta\wedge\tilde{\theta}^{2}-i\tilde{\theta}^{1}\wedge\tilde{\theta}^{3},$
$d\tilde{\theta}^{2}$ $=i\bigl{(}P(\zeta,\overline{\zeta})+P(\eta,\eta)\bigr{)}\theta\wedge\tilde{\theta}^{1}+iP(\zeta,\zeta)\theta\wedge\tilde{\theta}^{\overline{1}}$
$\quad-iP(\zeta,\eta)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}-i\tilde{\theta}^{2}\wedge\tilde{\theta}^{3},$
$d\tilde{\theta}^{3}$ $=-iP(\overline{\zeta},\eta)\theta\wedge\tilde{\theta}^{1}+iP(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{1}}$
$\quad+2iP(\zeta,\overline{\zeta})\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}+i\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{2}}.$
@@
$d\theta$ $=d\eta_{i}\wedge dx^{i}=i(\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{2}}+\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{1}}),$
$d\tilde{\theta}^{1}$ $=d\zeta_{i}\wedge dx^{i}=i\theta\wedge\tilde{\theta}^{2}-i\tilde{\theta}^{1}\wedge\tilde{\theta}^{3}.$
To compute $d\tilde{\theta}^{2}$, we note that
$d(\Gamma_{kjl}dx^{k})$ $=\frac{1}{2}R_{kmlj}dx^{k}\wedge dx^{m}$
$=\frac{1}{2}(P_{kl}g_{mj}-P_{ml}g_{kj}+P_{mj}g_{kl}-P_{kj}g_{ml})dx^{k}\wedge dx^{m}$
$=(P_{kl}g_{mj}+P_{mj}g_{kl})dx^{k}\wedge dx^{m}$
@@
$d\tilde{\theta}^{2}$ $=id\eta^{j}\wedge d\zeta_{j}-i\eta^{j}\zeta^{l}d(\Gamma_{kjl}dx^{k})$
$=id\eta^{j}\wedge d\zeta_{j}-i\eta^{j}\zeta^{l}(P_{kl}g_{mj}+P_{mj}g_{kl})dx^{k}\wedge dx^{m}$
$=i(\overline{\zeta^{j}}\tilde{\theta}^{2}-\zeta^{j}\tilde{\theta}^{\overline{2}})\wedge(\eta_{j}\tilde{\theta}^{2}-\zeta_{j}\tilde{\theta}^{3})$
$\quad-i\eta^{j}\zeta^{l}(P_{kl}g_{mj}+P_{mj}g_{kl})(\eta^{k}\theta+\overline{\zeta^{k}}\tilde{\theta}^{1}+\zeta^{k}\tilde{\theta}^{\overline{1}})\wedge(\eta^{m}\theta+\overline{\zeta^{m}}\tilde{\theta}^{1}+\zeta^{m}\tilde{\theta}^{\overline{1}})$
$=i\bigl{(}P(\zeta,\overline{\zeta})+P(\eta,\eta)\bigr{)}\theta\wedge\tilde{\theta}^{1}+iP(\zeta,\zeta)\theta\wedge\tilde{\theta}^{\overline{1}}-iP(\zeta,\eta)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}-i\tilde{\theta}^{2}\wedge\tilde{\theta}^{3}.$
The computation of $d\tilde{\theta}^{3}$ is similar, and we omit it. ∎
@@
$d\theta=i(\theta^{1}\wedge\theta^{\overline{2}}+\theta^{2}\wedge\theta^{\overline{1}})$
@@
$(h_{\alpha\overline{\beta}})=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$
@@
4.3. The Fefferman space for twistor CR manifolds
Let $\mathcal{C}=K_{M}^{\times}/\mathbb{R}_{+}$ be the Fefferman space over $M$. For simplicity, we assume that $\Sigma$ is globally oriented, and let $\theta=\eta_{i}dx^{i}$ be the contact form determined by a choice of representative metric $g\in[g]$ as above. We embed $\mathcal{C}$ into $K^{\times}_{M}$ as in (2.1) by $\theta$, and define the mapping
$\varpi\colon\widehat{\mathcal{C}}\longrightarrow\mathcal{C},\quad u\longmapsto(\theta\wedge\tilde{\theta}^{1}\wedge\tilde{\theta}^{2})_{u}$
by using our twisted coframe $(\tilde{\theta}^{1},\tilde{\theta}^{2})$ for $T^{1,0}M$. Since we have
$\delta_{\lambda}^{*}(\theta\wedge\tilde{\theta}^{1}\wedge\tilde{\theta}^{2})=\lambda^{2}(\theta\wedge\tilde{\theta}^{1}\wedge\tilde{\theta}^{2})$

### v4

@@ 4.1 An adapted frame on $\boldsymbol{\widehat{\mathcal{C}}}$
$\eta:={\rm i}\zeta\times\overline{\zeta}=2\operatorname{Re}\zeta\times\operatorname{Im}\zeta.$
@@
$\eta^{j}={\rm i}\varepsilon^{jkl}\zeta_{k}\overline{\zeta_{l}}.$
@@
$g(\eta,\eta)=1,\qquad$ $g(\eta,\zeta)=g\bigl(\eta,\overline{\zeta}\bigr)=0,$
$\eta\times\overline{\zeta}={\rm i}\overline{\zeta},\qquad$ $\zeta\times\eta={\rm i}\zeta.$
@@
Since the conditions $g(\zeta,\zeta)=0$, $g\bigl(\zeta,\overline{\zeta}\bigr)=1$ are preserved by the parallel transport of $\zeta$, we can define the horizontal lifts of tangent vectors on $\Sigma$ to $\widehat{\mathcal{C}}$. Let $\Gamma_{ij}{}^{k}$ be the Christoffel symbols of the Levi-Civita connection of $g$. Then, the horizontal lift of the tangent vector $\partial/\partial x^{i}$ is given by
@@
$\delta\zeta_{i}:={\rm d}\zeta_{i}-\Gamma_{ki}{}^{j}\zeta_{j}{\rm d}x^{k},\qquad\delta\overline{\zeta}_{i}:={\rm d}\overline{\zeta}_{i}-\Gamma_{ki}{}^{j}\overline{\zeta}_{j}{\rm d}x^{k},$
@@
Using these notations, we define a frame $\bigl(\tilde{T},\tilde{Z}_{1},\tilde{Z}_{2},\tilde{Z}_{\overline{1}},\tilde{Z}_{\overline{2}},K\bigr)$ for $\mathbb{C}T\widehat{\mathcal{C}}$ and its dual coframe $\bigl(\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\bigr)$ by
$\tilde{T}:=\eta^{i}\frac{\delta}{\delta x^{i}},\qquad\tilde{Z}_{1}:=\overline{\zeta^{i}}\frac{\delta}{\delta x^{i}},\qquad\tilde{Z}_{2}:=-{\rm i}\eta_{j}\frac{\partial}{\partial\zeta_{j}},\qquad K:={\rm i}\biggl(\zeta_{j}\frac{\partial}{\partial\zeta_{j}}-\overline{\zeta}_{j}\frac{\partial}{\partial\overline{\zeta}_{j}}\biggr)$
@@
$\theta:=\eta_{i}{\rm d}x^{i},\qquad\tilde{\theta}^{1}:=\zeta_{i}{\rm d}x^{i},\qquad\tilde{\theta}^{2}:={\rm i}\eta^{j}\delta\zeta_{j},\qquad\tilde{\theta}^{3}:=-\frac{{\rm i}}{2}\bigl(\overline{\zeta^{j}}\delta\zeta_{j}-\zeta^{j}\delta\overline{\zeta}_{j}\bigr)=-i\overline{\zeta^{j}}\delta\zeta_{j}.$
@@
We shall prove that this indeed gives a (co)frame which is ‘adapted’ to the CR structure on $M$.
@@
  • $(1)$
    The $1$-from $\theta$ descends to a contact form on $M$.
  • $(2)$
    The $1$-forms $\tilde{\theta}^{1}$, $\tilde{\theta}^{2}$ are horizontal with respect to $\widehat{\mathcal{C}}\to M$, and for any local section $^{s\colon M\to\widehat{\mathcal{C}}}$, the set of $1$-forms $\bigl(\theta,\theta^{1}=s^{*}\tilde{\theta}^{1},\theta^{2}=s^{*}\tilde{\theta}^{2},\theta^{\overline{1}}=\overline{\theta^{1}},\theta^{\overline{2}}=\overline{\theta^{2}}\bigr)$ gives an admissible coframe on $M$.
@@
(1) Since each coefficient $\eta_{i}$ is a $\mathrm{U}(1)$-invariant function, $\theta$ descends to a $1$-form on $M$. Moreover, by Proposition 3.3, $\theta$ annihilates $T^{1,0}M\oplus T^{0,1}M$ hence is a contact form on $M$. We note that the projection $T$ of $\tilde{T}$ to $M$ is well-defined since $\tilde{T}$ is $\mathrm{U}(1)$-invariant, and it gives the Reeb vector field of $\theta$. In fact, $T$ is represented as $T=\eta^{i}\bigl(\partial/\partial x^{i}\bigr)$ at $p$, and since ${\rm d}\varepsilon_{ijk}(x_{0})=0$, ${\rm d}g_{ij}(x_{0})=0$, we have ${\rm d}\eta_{i}(T)=0$, $\eta^{i}{\rm d}\eta_{i}=0$ at $p$ and hence $\theta(T)=\eta_{i}\eta^{i}=1$, $T\lrcorner\,{\rm d}\theta=T\lrcorner\,\bigl({\rm d}\eta_{i}\wedge{\rm d}x^{i}\bigr)=0$.
(2) Since $\tilde{\theta}^{1}(K)=\tilde{\theta}^{2}(K)=0$, these forms are horizontal. By Proposition 3.3, we see that $\theta^{1}$ and $\theta^{2}$ annihilate $T^{0,1}M$. Moreover, $\tilde{\theta}^{1}\bigl(\tilde{T}\bigr)=\tilde{\theta}^{2}\bigl(\tilde{T}\bigr)=0$ implies $\theta^{1}(T)=\theta^{2}(T)=0$. Thus, $\bigl(\theta,\theta^{1},\theta^{2},\theta^{\overline{1}},\theta^{\overline{2}}\bigr)$ is an admissible coframe. ∎
@@
$\delta_{\lambda}^{*}\tilde{\theta}^{1}=\lambda\tilde{\theta}^{1},\qquad\delta_{\lambda}^{*}\tilde{\theta}^{2}=\lambda\tilde{\theta}^{2}.$
Thus, $\bigl(\tilde{\theta}^{1},\tilde{\theta}^{2}\bigr)$ can be considered as a ‘twisted’ or ‘weighted’ coframe for $T^{1,0}M$.
We also note that if we replace $g$ with $\widehat{g}={\rm e}^{2\Upsilon}g$, $\Upsilon\in C^{\infty}(\Sigma)$, we have $\widehat{\eta}^{\,i}={\rm e}^{-\Upsilon}\eta^{i}$, $\widehat{\eta}_{i}={\rm e}^{\Upsilon}\eta_{i}$ and hence the associated contact form is rescaled as $\widehat{\theta}={\rm e}^{\Upsilon}\theta$.
4.2 The differentials of the adapted coframe
@@
${\rm d}x^{j}=\eta^{j}\theta+\overline{\zeta^{j}}\tilde{\theta}^{1}+\zeta^{j}\tilde{\theta}^{\overline{1}},\qquad{\rm d}\zeta_{j}=-{\rm i}\bigl(\eta_{j}\tilde{\theta}^{2}-\zeta_{j}\tilde{\theta}^{3}\bigr),\qquad{\rm d}\eta^{j}={\rm i}\bigl(\overline{\zeta^{j}}\tilde{\theta}^{2}-\zeta^{j}\tilde{\theta}^{\overline{2}}\bigr)$ (4.1)
@@
${\rm d}x^{j}\bigl(\tilde{T}\bigr)=\eta^{j},\qquad{\rm d}x^{j}\bigl(\tilde{Z}_{1}\bigr)=\overline{\zeta^{j}},\qquad{\rm d}x^{j}\bigl(\tilde{Z}_{2}\bigr)={\rm d}x^{j}(K)=0,$
@@
${\rm d}\zeta_{j}\bigl(\tilde{T}\bigr)={\rm d}\zeta_{j}\bigl(\tilde{Z}_{1}\bigr)={\rm d}\zeta_{j}\bigl(\tilde{Z}_{\overline{1}}\bigr)={\rm d}\zeta_{j}\bigl(\tilde{Z}_{\overline{2}}\bigr)=0,\qquad{\rm d}\zeta_{j}\bigl(\tilde{Z}_{2}\bigr)=-{\rm i}\eta_{j},\qquad{\rm d}\zeta_{j}(K)={\rm i}\zeta_{j},$
and obtain the formula for ${\rm d}\zeta_{j}$. Finally, using $d\varepsilon_{ijk}(x_{0})=0$, we compute as
${\rm d}\eta={\rm i}\,{\rm d}\zeta\times\overline{\zeta}+{\rm i}\zeta\times{\rm d}\overline{\zeta}=\bigl(\eta\tilde{\theta}^{2}-\zeta\tilde{\theta}^{3}\bigr)\times\overline{\zeta}-\zeta\times\bigl(\eta\tilde{\theta}^{\overline{2}}-\overline{\zeta}\tilde{\theta}^{3}\bigr)={\rm i}\overline{\zeta}\tilde{\theta}^{2}-{\rm i}\zeta\tilde{\theta}^{\overline{2}}$
@@
${\rm d}\theta={\rm i}\bigl(\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{2}}+\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{1}}\bigr),$
${\rm d}\tilde{\theta}^{1}={\rm i}\theta\wedge\tilde{\theta}^{2}-{\rm i}\tilde{\theta}^{1}\wedge\tilde{\theta}^{3},$
${\rm d}\tilde{\theta}^{2}={\rm i}\bigl(P\bigl(\zeta,\overline{\zeta}\bigr)+P(\eta,\eta)\bigr)\theta\wedge\tilde{\theta}^{1}+{\rm i}P(\zeta,\zeta)\theta\wedge\tilde{\theta}^{\overline{1}}-{\rm i}P(\zeta,\eta)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}-{\rm i}\tilde{\theta}^{2}\wedge\tilde{\theta}^{3},$
${\rm d}\tilde{\theta}^{3}=-{\rm i}P\bigl(\overline{\zeta},\eta\bigr)\theta\wedge\tilde{\theta}^{1}+{\rm i}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{1}}+2{\rm i}P\bigl(\zeta,\overline{\zeta}\bigr)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}+{\rm i}\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{2}}.$ (4.2)
@@
${\rm d}\theta={\rm d}\eta_{i}\wedge{\rm d}x^{i}={\rm i}\bigl(\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{2}}+\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{1}}\bigr),$
${\rm d}\tilde{\theta}^{1}={\rm d}\zeta_{i}\wedge{\rm d}x^{i}={\rm i}\theta\wedge\tilde{\theta}^{2}-{\rm i}\tilde{\theta}^{1}\wedge\tilde{\theta}^{3}.$
To compute ${\rm d}\tilde{\theta}^{2}$, we note that
${\rm d}\bigl(\Gamma_{kjl}{\rm d}x^{k}\bigr)$ $=\frac{1}{2}R_{kmlj}{\rm d}x^{k}\wedge{\rm d}x^{m}$
$=\frac{1}{2}(P_{kl}g_{mj}-P_{ml}g_{kj}+P_{mj}g_{kl}-P_{kj}g_{ml}){\rm d}x^{k}\wedge{\rm d}x^{m}$
$=(P_{kl}g_{mj}+P_{mj}g_{kl}){\rm d}x^{k}\wedge{\rm d}x^{m}$
@@
${\rm d}\tilde{\theta}^{2}={}$ ${\rm i}\,{\rm d}\eta^{j}\wedge{\rm d}\zeta_{j}-{\rm i}\eta^{j}\zeta^{l}{\rm d}\bigl(\Gamma_{kjl}{\rm d}x^{k}\bigr)$
$={}$ ${\rm i}\,{\rm d}\eta^{j}\wedge{\rm d}\zeta_{j}-{\rm i}\eta^{j}\zeta^{l}(P_{kl}g_{mj}+P_{mj}g_{kl}){\rm d}x^{k}\wedge{\rm d}x^{m}$
$={}$ ${\rm i}\bigl(\overline{\zeta^{j}}\tilde{\theta}^{2}-\zeta^{j}\tilde{\theta}^{\overline{2}}\bigr)\wedge\bigl(\eta_{j}\tilde{\theta}^{2}-\zeta_{j}\tilde{\theta}^{3}\bigr)$
${-}\,{\rm i}\eta^{j}\zeta^{l}(P_{kl}g_{mj}+P_{mj}g_{kl})\bigl(\eta^{k}\theta+\overline{\zeta^{k}}\tilde{\theta}^{1}+\zeta^{k}\tilde{\theta}^{\overline{1}}\bigr)\wedge\bigl(\eta^{m}\theta+\overline{\zeta^{m}}\tilde{\theta}^{1}+\zeta^{m}\tilde{\theta}^{\overline{1}}\bigr)$
$={}$ ${\rm i}\bigl(P\bigl(\zeta,\overline{\zeta}\bigr)+P(\eta,\eta)\bigr)\theta\wedge\tilde{\theta}^{1}+{\rm i}P(\zeta,\zeta)\theta\wedge\tilde{\theta}^{\overline{1}}-{\rm i}P(\zeta,\eta)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}-{\rm i}\tilde{\theta}^{2}\wedge\tilde{\theta}^{3}.$
The computation of ${\rm d}\tilde{\theta}^{3}$ is similar, and we omit it. ∎
@@
${\rm d}\theta={\rm i}\bigl(\theta^{1}\wedge\theta^{\overline{2}}+\theta^{2}\wedge\theta^{\overline{1}}\bigr)$
@@
$\bigl(h_{\alpha\overline{\beta}}\bigr)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$,
@@
4.3 The Fefferman space for twistor CR manifolds
Let $\mathcal{C}=K_{M}^{\times}/\mathbb{R}_{+}$ be the Fefferman space over $M$. For simplicity, we assume that $\Sigma$ is globally oriented, and let $\theta=\eta_{i}{\rm d}x^{i}$ be the contact form determined by a choice of representative metric $g\in[g]$ as above. We embed $\mathcal{C}$ into $K^{\times}_{M}$ as in (2.1) by $\theta$, and define the mapping
$\varpi\colon\ \widehat{\mathcal{C}}\longrightarrow\mathcal{C},\qquad u\longmapsto\bigl(\theta\wedge\tilde{\theta}^{1}\wedge\tilde{\theta}^{2}\bigr)_{u}$
by using our twisted coframe $\bigl(\tilde{\theta}^{1},\tilde{\theta}^{2}\bigr)$ for $T^{1,0}M$. Since we have
$\delta_{\lambda}^{*}\bigl(\theta\wedge\tilde{\theta}^{1}\wedge\tilde{\theta}^{2}\bigr)=\lambda^{2}\bigl(\theta\wedge\tilde{\theta}^{1}\wedge\tilde{\theta}^{2}\bigr)$